The rear wheel of a 150 kg tricycle accidentally ran over a nail. The circular tip of the nail has a diameter of 0.001 m. How much pressure did the wheel of the tricycle experience assuming that the entire weight of the tricycle is concentrated on the rear wheel that ran over the nail.

1
Expert's answer
2022-03-04T08:30:27-0500

Answer


Pressure is given by

"P=\frac{150*9.8*4}{3.14*0.001* 0.001}\\=187.26*10^{7}N/m"